# Diffusion approximation of the stationary distribution of a two-level single server queue

### v6

@@ 4.2 Preliminary lemmas
$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}1(L^{(r)}\leq\ell^{(r)}_{0})],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}1(L^{(r)}>\ell^{(r)}_{0})],$
$\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|L^{(r)}>\ell^{(r)}_{0}].$
@@
Functions $\eta^{(r)}_{i}(\theta)$ and $\zeta^{(r)}(\theta)$ play key roles in our analysis. Consider their properties. First, it is easy to see that they are infinitely many differentiable functions of $\theta$. Hence, twice differentiating the equations in (4.8), the following lemma can be obtained by their Taylor expansions around the origin. It is formally proved in [3, Lemma 5.8].
@@
Under the assumptions (3.a)–(3.d), we have the following expansions as $r\downarrow 0$.
$\eta_{i}^{(r)}(r\theta)=\lambda^{(r)}_{i}r\theta+\frac{1}{2}(\lambda^{(r)}_{i})^{3}(\sigma^{(r)}_{e,i})^{2}r^{2}\theta^{2}+o(r^{2}),\qquad i=1,2,$ (4.10)
$\zeta^{(r)}(r\theta)=-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}r^{2}\theta^{2}+o(r^{2}).$ (4.11)
@@
$\left|\eta^{(r)}_{i}(r\theta)(u_{1}\wedge 1/r)+\zeta^{(r)}(r\theta)(u_{2}\wedge 1/r)\right|\leq|\theta|\left(d_{e,i}(ru_{1}\wedge 1)+d_{s}(ru_{2}\wedge 1)\right).$ (4.12)
@@
because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$. These finiteness guarantees our computations for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$.
@@
Since $\lambda^{(r)}_{i}-c^{(r)}_{i}\mu^{(r)}=-b_{i}\mu^{(r)}r+o(r)$, it follows from Lemma 4.2, $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ and $\lambda_{i}=c_{i}\mu$ that
$\eta^{(r)}_{i}(r\theta_{i})+c^{(r)}_{i}\zeta^{(r)}(r\theta_{i})$ $=-b_{i}r^{2}\mu\theta_{i}+\frac{1}{2}c_{i}(\lambda_{i}^{2}\sigma_{e,i}^{2}+\mu^{2}\sigma_{s}^{2})r^{2}\mu\theta_{i}^{2}+o(r^{2})$
$=-\frac{1}{2}c_{i}\sigma_{i}^{2}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o(r^{2}).$
@@
$\qquad+c^{(r)}_{1}\left(1-\frac{1}{2}\mu^{2}\sigma_{s}^{2}r\theta_{1}\right)\mu r\theta_{1}\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})\right]+o(r^{2}).$ (4.17)
@@
$\widehat{T}^{(r)}_{e,i}\equiv T^{(r)}_{e,i}(0)\wedge 1/r,\quad i=1,2,\qquad\widehat{T}^{(r)}_{s}\equiv T^{(r)}_{s}(0-)\wedge 1/r,$
$\widehat{R}^{(r)}_{e}\equiv R^{(r)}_{e}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s}\equiv R^{(r)}_{s}(0-)\wedge 1/r.$
@@
$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$
$\quad=(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\qquad+e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}})1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$ (4.18)
@@
$\quad=\mathbb{E}^{(r)}_{s}\Big{[}e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0))1(L^{(r)}(0)=\ell^{(r)}_{0})-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0)))1(L^{(r)}_{1}(0)=\ell^{(r)}_{0}+1)\Big{]}$
@@
$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
@@
$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o(r^{2})\big{)}\psi^{(r)}_{i}(\theta_{i})$
$\quad-c_{1}\left(\mu r\theta_{1}+\mu^{3}\sigma_{s}^{2}r^{2}\theta_{1}^{2}/2+o(r^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})]=D^{(r)}(\theta_{1},\theta_{2}).$ (4.21)
This BAR is called an asymptotic BAR, and a starting point of our proof of Theorem 3.1. We will divide this BAR by $r^{2}$, and take its limit as $r\downarrow 0$, expanding the exponential functions in (4.3). To perform this expansion, we consider asymptotic behaviors on $\widehat{T}^{(r)}_{v}$ for $v=(e,1),(e,2),s$ and $\widehat{R}^{(r)}_{v^{\prime}}$ for $v^{\prime}=e,s$ as $r\downarrow 0$.
@@
Under the assumption (3.a), for $k=0,1,2$ as $r\downarrow 0$,
$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}1(R^{(r)}_{e}\in B)]\leq\alpha^{(r)}_{e}\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k+1}1(T^{(r)}_{e,i}\in B)],$ (4.22)
$(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[(R^{(r)}_{s})^{k}1(R^{(r)}_{s})\in B,L^{(r)}\geq 1)]$
$\qquad\leq\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{s})^{k+1}1(T^{(r)}_{s}\in B)],$ (4.23)
and, for $i=1,2$,
@@
$|\mathbb{E}[(\widehat{R}^{(r)}_{e})^{k}]-\mathbb{E}[(R^{(r)}_{e})^{k}]|=\mathbb{E}[(R^{(r)}_{v})^{k}1(R^{(r)}_{v}>1/r)]=o(r^{2-k}),\qquad v=e,s,$ (4.25)
$\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{k}]=\mathbb{E}^{(r)}_{s}[(R^{(r)}_{e}(0))^{k}]+o(r^{3-k}),$ (4.26)
$\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{k}]=\mathbb{E}^{(r)}_{e}[(R^{(r)}_{s}(0-))^{k}]+o(r^{3-k}).$ (4.27)
In particular, $\{(R^{(r)}_{e})^{k};r\in(0,1]\}$ and $\{(R^{(r)}_{s})^{k}1(L^{(r)}(0)\geq 1);r\in(0,1]\}$ are uniformly integrable for $k=1,2$.
This lemma is easily proved using the BAR (4.1), but its proof is lengthy, so we defer it to Appendix A.3. Because of the finiteness of the third moments in the assumption (3.a), the following corollary is immediate from Lemma 4.4.
Corollary 4.3.
For each $k=1,2$, $\mathbb{E}_{e}[(\widehat{R}^{(r)}_{s})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are uniformly bounded for $r\in(0,1]$.
Note that the weak limit of $\nu^{(r)}$ should be obtained from (4.3). However, we will see that it requires the relation between the limits of $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, which is not easy to get from (4.3). To relax this difficulty, we prepare two lemmas below.
@@
Under the assumption (3.a), we have, for $k=1,2,3$ and as $r\downarrow 0$,
@@
$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]\Big{)},$ (4.28)
@@
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)>\ell^{(r)}_{0}]\Big{)},$ (4.29)
@@
$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,2})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)\geq\ell^{(r)}_{0}]\Big{)},$ (4.31)
@@
For $k\geq 1$, we first prove (4.5). We apply $f({\boldsymbol{x}})=(x_{3}\wedge 1/r)^{k}1(x_{1}\leq\ell^{(r)}_{0})$ to (4.1), then we have (4.5) because
@@
Similarly, (4.5), (4.5) and (4.5) are obtained from (4.1) by applying $f({\boldsymbol{x}})=(x_{3}\wedge 1/r)^{k}1(x_{1}>\ell^{(r)}_{0})$, $f({\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(x_{1}\leq\ell^{(r)}_{0})$ and $f({\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(x_{1}>\ell^{(r)}_{0})$, respectively. ∎
Define, for $i=1,2$,
$\Delta^{(r)}_{i}$ $=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
@@
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-r\mu b_{1}\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.33)
@@
$\hskip 43.05542pt-\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\lambda^{(r)}_{1}\mathbb{E}_{s}^{(r)}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0}]\Big{)}+o(r),$
because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1 and $\mathbb{P}(R^{(r)}_{e}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ and $\mathbb{P}(R^{(r)}_{s}\geq 1/r,L^{(r)}>\ell^{(r)}_{0})=o(r^{2})$ by (4.25). This proves (4.33). Similarly, (4.34) is obtained from (4.5) and (4.5), ∎
@@
4.3 Proof of Theorem 3.1
We first prove (i). We are now ready to take the limit of the BAR of (4.3) as $r\downarrow 0$. For this, we will use the following elementary expansions.

### v7

@@ 4.2 Asymptotic BAR for the pre-limit process $X^{(r)}(\cdot)$
$\varphi^{(r)}_{1}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}1(L^{(r)}\leq\ell^{(r)}_{0})],\qquad\varphi^{(r)}_{2}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}1(\ell^{(r)}_{0}<L^{(r)})],$
$\varphi^{(r)}_{1|cd}(\theta_{1})=\mathbb{E}[e^{\theta_{1}rL^{(r)}}|L^{(r)}\leq\ell^{(r)}_{0}],\qquad\varphi^{(r)}_{2|cd}(\theta_{2})=\mathbb{E}[e^{\theta_{2}rL^{(r)}}|\ell^{(r)}_{0}<L^{(r)}].$
@@
Functions $\eta^{(r)}_{i}(\theta)$ and $\zeta^{(r)}(\theta)$ play key roles in our analysis. Consider their properties. First, it is easy to see that they are infinitely many differentiable functions of $\theta$. Hence, twice differentiating the equations in (4.8), the following lemma can be obtained by their Taylor expansions around the origin. It is formally proved in [4, Lemma 5.8].
@@
Under the assumptions (3.a)–(3.d), we have the following expansions as $r\downarrow 0$ for each fixed $\theta\in\mathbb{R}$.
$\eta_{i}^{(r)}(r\theta)=\lambda^{(r)}_{i}r\theta+\frac{1}{2}(\lambda^{(r)}_{i})^{3}(\sigma^{(r)}_{e,i})^{2}r^{2}\theta^{2}+o((\theta r)^{2}),\qquad i=1,2,$ (4.10)
$\zeta^{(r)}(r\theta)=-\mu^{(r)}r\theta+\frac{1}{2}(\mu^{(r)})^{3}(\sigma^{(r)}_{s})^{2}r^{2}\theta^{2}+o((\theta r)^{2}).$ (4.11)
@@
$\left|\eta^{(r)}_{i}(r\theta)(u_{1}\wedge 1/r)\right|\leq|r\theta|d_{e,i}(u_{1}\wedge 1/r),\qquad i=1,2,$ (4.12)
$\left|\zeta^{(r)}(r\theta)(u_{2}\wedge 1/r)\right|\leq|r\theta|d_{s}(u_{2}\wedge 1/r).$ (4.13)
Remark 4.2.
In (4.10) and (4.11), $o((\theta r)^{2})$ can be replaced by $o(r^{2})$ because $\theta$ is fixed, but $o((\theta r)^{2})$ is convenient for our computation because we need to divide both sides of them by $\theta$ in some cases. We will take this convention throughout the paper as long as it is possibe.
@@
because $r\theta_{1}L^{(r)}1(L^{(r)}\leq\ell^{(r)}_{0})\leq\theta_{1}(\ell+o(r))$ for $\theta_{1}>0$. These finiteness guarantees our computations for $\theta_{1}\in\mathbb{R}$ and $\theta_{2}\leq 0$. We also note the following fact.
Corollary 4.3.
For $\theta_{1}\in\mathbb{R}_{+}$ and $\theta_{2}\leq 0$
$\lim_{r\downarrow 0}\left|\psi^{(r)}_{i}(\theta_{i})-\varphi^{(r)}_{i}(\theta_{i}))\right|=0,\quad i=1,2.$ (4.17)
Proof.
By Corollary 4.2, we have
This proves (4.17). ∎
@@
Since $\lambda^{(r)}_{i}-c^{(r)}_{i}\mu^{(r)}=-b_{i}\mu^{(r)}r+o(r)$, it follows from Lemma 4.2, $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ and $\lambda_{i}=c_{i}\mu$ that, each fixed $\theta_{i}\in\mathbb{R}$,
$\eta^{(r)}_{i}(r\theta_{i})+c^{(r)}_{i}\zeta^{(r)}(r\theta_{i})$ $=-b_{i}r^{2}\mu\theta_{i}+\frac{1}{2}c_{i}(\lambda_{i}^{2}\sigma_{e,i}^{2}+\mu^{2}\sigma_{s}^{2})r^{2}\mu\theta_{i}^{2}+o((\theta_{i}r)^{2})$
$=-\frac{1}{2}c_{i}\sigma_{i}^{2}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o((\theta_{i}r)^{2}),\qquad r\downarrow 0.$
@@
$\quad+c^{(r)}_{1}\left(1-\frac{1}{2}\mu^{2}\sigma_{s}^{2}r\theta_{1}\right)\mu r\theta_{1}\mathbb{E}\left[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})\right]+o((\theta_{1}r)^{2})+o((\theta_{2}r)^{2}).$ (4.19)
@@
$\widehat{T}^{(r)}_{e,i}\equiv T^{(r)}_{e,i}\wedge 1/r,\quad i=1,2,\qquad\widehat{T}^{(r)}_{s}\equiv T^{(r)}_{s}\wedge 1/r,$ (4.20)
$\widehat{R}^{(r)}_{e}\equiv R^{(r)}_{e}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s}\equiv R^{(r)}_{s}(0)\wedge 1/r,\qquad\widehat{R}^{(r)}_{s-}\equiv R^{(r)}_{s}(0-)\wedge 1/r.$ (4.21)
@@
$\quad=e^{r\theta_{2}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\hskip 64.58313pt-e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$
$\quad=(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\qquad+e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0}))],$ (4.22)
@@
$\quad=\mathbb{E}^{(r)}_{s}\Big{[}e^{r\theta_{1}\ell^{(r)}_{0}}g^{(r)}_{1,r\theta_{1}}(R^{(r)}(0))1(L^{(r)}(0)=\ell^{(r)}_{0})-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}g^{(r)}_{2,r\theta_{2}}(R^{(r)}_{1}(0)))1(L^{(r)}_{1}(0-)=\ell^{(r)}_{0}+1)\Big{]}$
@@
$D^{(r)}(\theta_{1},\theta_{2})=\alpha^{(r)}_{e}(e^{r\theta_{2}\ell^{(r)}_{0}}-e^{r\theta_{1}\ell^{(r)}_{0}})\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}e^{r\theta_{1}\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
@@
$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta_{i})\mu\theta_{i}r^{2}+o((\theta r)^{2})\big{)}\psi^{(r)}_{i}(\theta_{i})$
$\quad-c_{1}\left(\mu r\theta_{1}-\mu^{3}\sigma_{s}^{2}r^{2}\theta_{1}^{2}/2+o((\theta_{1}r)^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta_{1}}(R^{(r)})]=D^{(r)}(\theta_{1},\theta_{2}).$ (4.25)
This BAR is called an asymptotic BAR, and a starting point of our proof of Theorem 3.1.
4.3 Expansion of the asymptotic BAR
We aim to derive the limits of $\psi^{(r)}_{1}(\theta_{1})$ and $\psi^{(r)}_{2}(\theta_{2})$ from the BAR (4.3). For this, we will divide it for either $\theta_{2}=0$ or $\theta_{1}=0$ by $r^{2}$, and take its limit as $r\downarrow 0$, expanding the exponential functions in (4.3). We perform this computations in three steps. Recall that (3.a)–(3.e) are assumed but (3.f) is not assumed throughout Section 4.
(Step 1)   In this step, we show that the moments of $\widehat{T}^{(r)}_{v}$ for $v=(e,1),(e,2),s$ and $\widehat{R}^{(r)}_{v^{\prime}}$ for $v^{\prime}=e,s$ are well behaved as $r\downarrow 0$. This is also examined for $\widehat{R}^{(r)}_{s-}1(L^{(r)}=\ell^{(r)}_{0})$ and $\widehat{R}^{(r)}_{e}1(L^{(r)}=\ell^{(r)}_{0-})$ under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$, respectively. These are basis for the expansion of the BAR. We then introduce the set of key quantities, $\Delta^{(r)}_{i}$ for $i=1,2$.
@@ 4.3 Expansion of the asymptotic BAR
Under the assumption (3.a), for $k=0,1,2$ and $r\in(0,1]$,
$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}]\leq\alpha^{(r)}_{e}\max_{i=1,2}\mathbb{E}[(T^{(r)}_{e,i})^{k+1}],$ (4.26)
$(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[(R^{(r)}_{s})^{k}]\leq\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{s})^{k+1}]+(k+1)(\min_{i=1,2}c^{(r)}_{i})\mathbb{E}[T_{s}^{k}],$ (4.27)
and, for $i=1,2$, as $r\downarrow 0$
@@
$|\mathbb{E}[(\widehat{R}^{(r)}_{v})^{k}]-\mathbb{E}[(R^{(r)}_{v})^{k}]|=\mathbb{E}[(R^{(r)}_{v})^{k}1(R^{(r)}_{v}>1/r)]=o(r^{2-k}),\qquad v=e,s.$ (4.29)
In particular, $\{(R^{(r)}_{e})^{k};r\in(0,1]\}$ and $\{(R^{(r)}_{s})^{k}1(L^{(r)}(0)\geq 1);r\in(0,1]\}$ are uniformly bounded for $k=1,2$ by (3.a).
This lemma is easily proved using the BAR (4.1), but its proof is lengthy, so we defer it to Appendix A.3. We prepare one more lemma and its corollary, which are helpful for the expansion of the BAR.
@@
Under the assumption (3.a), we have, for $k=1,2,3$ as $r\downarrow 0$,
@@
$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]\Big{)},$ (4.30)
@@
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{s})^{k}]\mathbb{P}^{(r)}_{s}[L^{(r)}(0)>\ell^{(r)}_{0}]\Big{)},$ (4.31)
@@
$\quad=\alpha^{(r)}_{e}\Big{(}-\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]+\mathbb{E}[(\widehat{T}^{(r)}_{e,2})^{k}]\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)\geq\ell^{(r)}_{0}]\Big{)}.$ (4.33)
@@
For $k\geq 1$, we first prove (4.5). We apply $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{0})$ to (4.1), then we have (4.5) because
@@
Similarly, (4.5), (4.5) and (4.5) are obtained from (4.1) by applying $f(z,{\boldsymbol{x}})=(x_{2}\wedge 1/r)^{k}1(z>\ell^{(r)}_{0})$, $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z\leq\ell^{(r)}_{0})$ and $f(z,{\boldsymbol{x}})=(x_{1}\wedge 1/r)^{k}1(z>\ell^{(r)}_{0})$, respectively. ∎
The following corollary is immediate from (4.5), (4.5), (3.a) and Lemma 4.4.
Corollary 4.4.
For each $k=1,2,3$, $\mathbb{E}_{e}[(\widehat{R}^{(r)}_{s-})^{k}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}_{s}[(\widehat{R}^{(r)}_{e})^{k}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are uniformly bounded for $r\in(0,1]$.
Note that the limits of $\psi^{(r)}_{1}(\theta_{1})$ and $\psi^{(r)}_{2}(\theta_{2})$ should be obtained from (4.3). However, we will see that it requires the relation between the limits of $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, which is not easy to get from (4.3). To consider this problem, we introduce the following quantities. Define, for $i=1,2$,
$\Delta^{(r)}_{i}$ $=\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
@@
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-rb_{1}\mu\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.35)
@@
$\hskip 43.05542pt-\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\lambda^{(r)}_{1}\mathbb{E}_{s}^{(r)}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0}]\Big{)}+o(r),$
because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1 and $\mathbb{P}(R^{(r)}_{e}\geq 1/r)=o(r^{2})$ and $\mathbb{P}(R^{(r)}_{s}\geq 1/r)=o(r^{2})$ by (4.29). This proves (4.35). Similarly, (4.36) is obtained from (4.5) and (4.5), ∎
@@
(Step 2)   In this step, we expand $D^{(r)}(\theta,0)$ and $D^{(r)}(0,\theta)$ concerning $r$ as $r\downarrow 0$. We will see that $\Delta^{(r)}_{i}$’s describe the expansion terms of $D^{(r)}(\theta,0)$ of order $r$. We here will introduce ${\mathcal{E}}^{(r)}_{i}$’s for describing their expansion terms of order $r^{2}$.
We start with the following elementary expansions. Recall Remark 4.2 for our convention of the order terms.